# Intuitionism with Truth Tables: A Decision Procedure for IPL Based on RNmatrices

### v2

@@ Abstract
This paper provides a new decision procedure for intuitionistic propositional logic (IPL) based on a 3-valued restricted non-deterministic matrix and proves its soundness and completeness.
@@ 1 Introduction
A famous result, proved first by K. Gödel in $1932$, says that it is impossible to characterize intuitionistic propositional logic (IPL) via a finite-valued (deterministic) matrix [1]. In other words, Gödel proved that there is no effective decision procedure based on finite-valued ordinary truth-tables for IPL.
Later, in $1940$, a maybe less famous but equally important result was proved by Dugundji, who adapted Gödel’s argument to establish that the same is also true about the family of modal logics between S1 and S5 [1].
To overcome Dugundji’s result, J. Kearns [6] introduced a $4$-valued Nmatrix (non-deterministic matrix) for modal logics KT, S4, and S5 and restricted the set of valuations to a subset called “level valuations”. He proved that this restricted Nmatrix (hereafter called RNmatrix [3]) constitutes sound and complete semantics for these modal logics.
Unfortunately, in principle, Kearns’s level valuations fail to provide an effective decision procedure for any of the modal logics. More than that, until recently, it was unknown whether obtaining such a decision procedure from level valuations would be possible.
@@
L. Grätz recently solved this problem for some modal systems [4]. He refined Kearn’s original RNmatrix to obtain a decidable $3$-valued RNmatrix for modal logics KT and S4. He noticed that, with an appropriate notion of partial valuation for level semantics (i.e., preserving analyticity and co-analyticity), the truth tables could be obtained by removing any spurious row eventually produced in the valuation process.
In this paper, we will try to overcome the original problem posed by Gödel on the characterizability of IPL via finite-valued matrix. We will propose a decidable 3-valued RNmatrix for IPL combining Grätz’s RNmatrix for S4 and Gödel-McKinsey-Tarski’s box translation.
2 Nmatrices and RNmatrices
Along this paper, all the propositional languages will be defined over a given denumerable set $\mathcal{P}=\{p_{0},p_{1},p_{2},\ldots\}$ of propositional variables.
Definition 2.1 (Nmatrix).
@@ 2 Nmatrices and RNmatrices
Let $\Theta$ be a signature and $For(\Theta)$ the propositional language over signature $\Theta$ generated by the set $\mathcal{P}$ of propositional variables. An Nmatrix for $\mathcal{L}$ over $\Theta$ is a tuple $\mathcal{M}=\langle\mathcal{V},\mathcal{D},\mathcal{O}\rangle$ where
@@
    $\mathcal{V}$ is the (non-empty) set of truth values;
@@
    $\mathcal{D}\subset\mathcal{V}$ is the (non-empty) set of designated values;
@@
    $\mathcal{O}$ is a function assigning, to each $n$-ary connective ##\## of $\Theta$, a (non-deterministic) truth function $\mathcal{O}(\#):\mathcal{V}^{m}\rightarrow(\wp(\mathcal{V}))\backslash\emptyset$.
A valuation over an Nmatrix $\mathcal{M}$ is a function $v:For(\Theta)\to\mathcal{V}$ such that, for every $n$-ary connective #normal-#\## of $\Theta$ and for every $\varphi_{1},\ldots,\varphi_{n}\in For(\Theta)$, $v(\#(\varphi_{1},\ldots,\varphi_{n}))\in\mathcal{O}(\#)(v(\varphi_{1}),\ldots,v(\varphi_{n}))$. The set of valuations over $\mathcal{M}$ is denoted by $Val(\mathcal{M})$.
Definition 2.2 (RNmatrix).
Let $\Theta$ be a signature. An RNmatrix [3] over $\Theta$ is a pair $\mathcal{R}=\langle\mathcal{M},\mathcal{F}\rangle$ where $\mathcal{M}$ is an Nmatrix over $\Theta$ and $\mathcal{F}\subseteq Val(\mathcal{M})$.
Definition 2.3 (Consequence relation in an (R)Nmatrix).
Let $\mathcal{M}$ ($\mathcal{R}$, resp.) be an Nmatrix (an RNmatrix, resp.) over $\Theta$, and $\Gamma\cup\{\varphi\}\subseteq For(\Theta)$. We say that $\varphi$ is a semantical consequence of $\Gamma$ w.r.t. $\mathcal{M}$ (w.r.t. $\mathcal{R}$, resp.), denoted by $\Gamma\models_{\mathcal{M}}\varphi$ (denoted by $\Gamma\models_{\mathcal{R}}\varphi$, resp.) if, for every $v\in Val(\mathcal{M})$ (for every $v\in\mathcal{F}$, resp.): $v(\varphi)\in\mathcal{D}$ whenever $v(\gamma)\in\mathcal{D}$ for every $\gamma\in\Gamma$. We write $\models_{\mathcal{M}}\varphi$ and $\models_{\mathcal{R}}\varphi$ instead of $\emptyset\models_{\mathcal{M}}\varphi$ and $\emptyset\models_{\mathcal{R}}\varphi$.
$\neg^{\mathcal{A}_{S4}}$
@@ 2 $\{0\}$$\{1,2\}$$\{1,2\}$
Example 2.4 ($\mathcal{M}_{S4}$).
Let $\Sigma_{b}=\{\{\neg,\Box\},\{\rightarrow\}\}$ and Σ={{¬,},{,,}}normal-$\Sigma=\{\{\neg,\Box\},\{\rightarrow,\lor,\land\}\}$ be, respectively, the basic signature and the signature for modal logic S4. Given the set $\mathcal{P}$ of propositional variables, let $For(\Sigma_{b})$ and $For(\Sigma)$ be the corresponding propositional languages. Grätz’s Nmatrix over $\Sigma_{b}$ is the tuple $\mathcal{M}_{S4}=\langle\mathcal{V},\mathcal{D},\mathcal{O}\rangle$, where $\mathcal{V}=\{0,1,2\}$, $\mathcal{D}=\{1,2\}$, $\mathcal{O}=\{\neg^{\mathcal{A}_{S4}},\Box^{\mathcal{A}_{S4}},\rightarrow^{\mathcal{A}_{S4}}\}$ and each $c\in\mathcal{O}$ is defined in Table 1.
$p$$\Box p$$\Box\Box p$$\Box p\rightarrow\Box\Box p$000121001222212
(a) Valuations in $\mathcal{M}_{S4}$ for $\Box p\rightarrow\Box\Box p$
Figure 1: Each tree level corresponds to the set of values resulting from the multioperation applied to the respective subformula (at left) in that tree. At the top are the values that the atomic formula can receive, and in the leaves are the values of the evaluated formula.
By setting $\varphi\lor\psi:=\neg\varphi\to\psi$ and $\varphi\land\psi:=\neg(\varphi\to\neg\psi)$, respectively, the multioperators $\lor^{\mathcal{A}_{S4}}$ and $\land^{\mathcal{A}_{S4}}$ are defined as in Table 1.
Valuations in an Nmatrix are typically represented as tables. However, if one focuses more on the computational aspects of these valuations, adopting the more flexible data structure of binary trees can be advantageous. In Figure 1, we show how to compute the valuation of two examples.
3 Kearns’s level valuations
Regarding semantics for modal logic, Kripke-style semantics are the usual choice. However, there are other attractive alternatives. In the 1980s, Kearns proposed an RNmatrix-based semantics for modal logics KT, S4 and S5. He introduced the level valuations as the set $\mathcal{F}$ of valuations to be used.
[Kearn’s account] is simpler than [Kripke’s account] in virtue of having dispensed with possible worlds and their relations. I also think that my account is philosophically preferable to the standard account for having done this. For I do not think there are such things as possible worlds, or even that they constitute a useful fiction. [6]
We will define the level valuations for the modal logic S4.
Figure 2: Assume that $Val(\mathcal{M}_{S4})=\{v_{0},v_{1},\ldots,v_{m},v_{n},v_{k},\ldots\}$. We remove a valuation $v$ in a level $\mathcal{L}_{i+1}$ iff there is some wf $\alpha$ such that $v(\alpha)=1$ and $w(\alpha)\in\mathcal{D}$ for every $w\in\mathcal{L}_{i}$.

### v4

# A new decision method for Intuitionistic Logic by 3-valued non-deterministic truth-tables (pre-print version)

@@ Abstract
Kurt Gödel proved that it is not possible to characterize Intuitionistic Propositional Logic ($\mathbf{IPL}$) by means of finite and deterministic truth-tables. After extending the same result with respect to non-deterministic matrices, we provide a semantical characterization of $\mathbf{IPL}$ by means of a $3$-valued non-deterministic matrix with a restricted set of valuations. This structure allows to define an algorithm to delete unsound rows from the non-deterministic truth-tables generated for each formula, which constitutes a new and very simple decision procedure for $\mathbf{IPL}$. This method can be seen as truth-tables in a broader sense, and a way to overcome Gödel’s limiting result.
MSC class: 03B20; 03B45; 03B25; 03B35; 68V15
@@ 1 Introduction
In 1932, Kurt Gödel established the uncharacterizability of intuitionistic propositional logic (IPL) with respect to finite-valued deterministic matrices [10]. He demonstrated that any finite deterministic matrix, strongly enough to validate every intuitionistic tautology, will also validate at least one intuitionistically invalid proposition. Later, in $1940$, James Dugundji proved a similarly significant result by adapting Gödel’s argument to show that the same holds true for modal logic S5 and all of its subsystems [8].
These limitations led to new interpretations for modal and intuitionistic systems, such as Saul Kripke’s relational semantics, where the notion of possible world plays a central role. The tradition of using possible worlds to explain modalities dates back to Leibniz and is at the core of philosophical debates: Are possible worlds real or just helpful fiction? John Kearns challenged this tradition by defending that they are neither one nor the other [19, p. 86]. To support his position, he proposed, in $1981$, a four-valued semantics for the modal systems KT, S4, and S5 where the truth-value of a formula depends solely on the truth assignments of its subformulas. In Kearns’ approach, the meaning of certain connectives is not entirely determined by their matrices, as some inputs can yield more than one output. This type of semantics was later formalized under the notion of non-deterministic matrices (Nmatrices, in short) [2]. To address this indeterminacy, instead of possible worlds, Kearns introduced a method for restricting the set of valuations, which was later referred to more generally as level valuations (see Section 4). In this approach, not all valuations in the Nmatrix are permitted. Instead, only a subset that meets specific conditions to ensure the validity of the necessitation rule is allowed; this subset is precisely the set of level valuations. Nmatrices with a restricted set of valuations lead to the notion of Restricted non-deterministic matrices or Restricted Nmatrices (RNmatrices, in short), central in the present paper.
Unlike Kripke semantics, however, Kearns’s level valuations do not draw a straightforward decision procedure for those logics. Until recently, it was unclear whether obtaining such a decision procedure from level valuations would be possible.
@@
In $2021$, Lukas Grätz solved this problem for some modal systems (see [12]). He built upon Kearns’s original RNmatrix $3$-valued RNmatrices for KT and S4 with two decidable notions of partial level valuation, one for KT and other for S4 and show that both are analytic and co-analytic regarding the level valuations of the RNmatrix for the respective modal systems. To obtain the truth-table, one discards (by means of a simple algorithm) any valuation that does not fit the corresponding definition. Ultimately, the remaining partial valuations suffice to decide the formula’s validity. Analyticity and co-analyticity ensure that partial level valuation semantics (which constitute the rows of the truth-tables) coincide with the corresponding RNmatrix semantics, which characterizes each logic.
Constructing non-deterministic (i.e., row-branching) truth-tables in a finite-valued Nmatrix can be seen as a generalization of the standard (deterministic) truth-tables method for deciding validity in a given logic. Indeed, non-deterministic matrices were formally introduced by A. Avron and I. Lev in [2] with the aim of overcoming Gödel-Dugundji’s uncharacterizability results for some logics. The main purpose of Nmatrices was to obtain a decision procedure, analogous to finite-valued truth-tables, for logics that do not admit such a characterization. It is worth noting that Nmatrices were already considered in the literature before these authors. For instance, N. Rescher already defined a non-deterministic matrix semantics in 1962, under the name of quasi-truth-functional systems (see [28]). Afterwards, and among other authors, this notion was independently proposed by Y. Ivlev for studying non-normal modal logics, starting in the 1970s (see for instance [14, 15, 16, 17]), as well as J. Kearns (with restricted valuations).
In spite of its greater expressive power with respect to standard (deterministic) matrices, the generalization of the notion of truth-tables by means of Nmatrices has also limitations: some logics, such as da Costa’s paraconsistent logic $C_{1}$, cannot be characterized by a single finite Nmatrix, as proved by Avron in [1, Theorem 11]. The same is true for KT and S4, as proved by Grätz in [12]. Moreover, in Theorem 3.5 below we will show that $\mathbf{IPL}$ also cannot be characterized by a single finite Nmatrix, extending so Gödel’s theorem from $1932$. These examples show that Gödel-Dugundji’s results hold, for some logics, in a stronger sense, that is, w.r.t. Nmatrices. The key to overcoming this limitation is to consider some valuations in a suitable Nmatrix. This kind of semantical structures was formally introduced in [7], under the name of restricted Nmatrices, or RNmatrices, showing that each da Costa’s logics $C_{n}$ (for $n\geq 1$) can be characterized by a single (decidable) $(n+2)$-valued RNmatrix; in particular, validity in $C_{1}$ can be decided by a $3$-valued RNmatrix by means of branching truth-tables (plus a simple criterion to delete inadequate rows). Analogously, Grätz decidable $3$-valued RNmatrices produce a decision procedure for KT and S4 by means of $3$-valued branching truth-tables (together with a simple method to delete unsound rows).
Finite-valued non-deterministic truth-tables, together with an algorithm to delete inadequate rows, is a decision procedure that can be seen as truth-tables in a broader sense, and is a way to overcome Gödel-Dugundji’s limiting results, even for more complicated logics such as KT, S4, $C_{1}$ and $\mathbf{IPL}$. The aim of this paper is, precisely, to obtain a new decision procedure for $\mathbf{IPL}$ based on this kind of generalized truth-tables.
Recall that, in $1933$, Gödel proposed an interpretation of $\mathbf{IPL}$ into S4 by means of a translation mapping $t$ between their languages [11]. In $1948$, Tarski and McKinsey prove that Gödel’s mapping $t$ constitutes a faithful interpretation, that is: $\alpha$ is a theorem of $\mathbf{IPL}$ iff its translation $t(\alpha)$ is a theorem of S4 [25]. Since Gödel translation $t$ is computable, then, by composing it with any algorithm for deciding S4, a decision procedure for $\mathbf{IPL}$ is immediately obtained, defined by using a translation (from $\mathbf{IPL}$ to S4). In particular, by composing Gödel translation $t$ with Grätz’s algorithm for S4, a decision procedure is obtained for $\mathbf{IPL}$, based on an RNmatrix for S4 (as well as the translation $t$). The definition of a pure RNmatrix for $\mathbf{IPL}$ (i.e., without depending on Gödel and Grätz’s constructions), which is also decidable by truth-tables, as discussed above, is a challenge that naturally arises.
A pure RNmatrix for $\mathbf{IPL}$ with such features is desirable for several reasons. First, it can lead to improvements in efficiency. As will become clear later, the computational cost of producing a truth-table is directly related to the number of subformulas, and the translated formulas are twice as large as the original intuitionistic formulas. The second reason is that a pure RNmatrix can explain the actual meaning of intuitionistic operators. As it will be argued below, the RNmatrix for $\mathbf{IPL}$ proposed here has significant differences when compared to the RNmatrix for S4, and those differences are motivated by philosophical considerations on the nature of $\mathbf{IPL}$ itself.
In this paper, a pure (and decidable by non-deterministic truth-tables) RNmatrix for $\mathbf{IPL}$ will be constructed by abstracting the composition of Gödel’s translation algorithm and Grätz’s decision procedure for S4. This abstraction will provide a direct way to define the RNmatrix for $\mathbf{IPL}$. Hence, a $3$-valued Nmatrix, together with suitable notions of level valuations and partial level valuations (different, in nature, to the ones for modal logics) is introduced, and the soundness and completeness of the method are proved without relying on the results of Gödel [11] and Grätz [12].
The paper is organized as follows. Section 2 introduces the central notions that will be used, such as the notion of non-deterministic matrices, restricted non-deterministic matrices, and their valuation functions. Section 3 extends Gödel’s uncharacterizability theorem for $\mathbf{IPL}$ to the non-deterministic case. Section 4 presents Kearns’s Nmatrices, and the notion of level valuation for S4 is detailed. Section 5 presents Grätz’s RNmatrix for S4. Section 7 expands the RNmatrix for S4 with conjunction and disjunction and introduces Gödel’s box-translation. Section 8 introduces the RNmatrix for $\mathbf{IPL}$ and proves its soundness and completeness regarding a new notion of level valuation for $\mathbf{IPL}$. Section 9 presents the truth-table method based on the RNmatrix, and some examples are discussed. Section 10 briefly considers some aspects of the computational complexity of the method. Finally, Section 11 discusses some related works and possibilities of future research.
2 Preliminaries: Logics, Nmatrices and RNmatrices
Along this paper, a propositional signature is a denumerable family $\Theta=(\Theta_{n})_{n\geq 0}$ of pairwise disjoint sets. Elements in $\Theta_{n}$ are called $n$-ary connectives. An algebra over $\Theta$ is a pair $\mathcal{A}=\langle A,\mathcal{O}\rangle$ such that $A$ is a non-empty set called the universe or domain of $\mathcal{A}$ and, $\mathcal{O}$ is a function assigning, to each $n$-ary connective #\# of $\Theta$, a function $\mathcal{O}(\#):A^{n}\rightarrow A$. The (absolutely) free algebra over $\Theta$ generated by a denumerable set $\mathcal{P}=\{p_{0},p_{1},\ldots\}$ of propositional variables is called the algebra of formulas over $\Theta$, and will be denoted by $For(\Theta)$.
A logic is a pair ${\bf L}=\langle For,\vdash_{\bf L}\rangle$ where $For$ is the set of formulas and $\vdash_{\bf L}$ is the consequence relation. A logic L is said to be Tarskian if it satisfies the following: (1) if $\varphi\in\Gamma$ then $\Gamma\vdash_{\bf L}\varphi$; (2) if $\Gamma\vdash_{\bf L}\varphi$ and $\Gamma\subseteq\Delta$ then $\Delta\vdash_{\bf L}\varphi$; and (3) if $\Delta\vdash_{\bf L}\varphi$ and $\Gamma\vdash_{\bf L}\psi$ for every $\psi\in\Delta$ then $\Gamma\vdash_{\bf L}\varphi$. A Tarskian logic L is finitary if it satisfies: (4) if $\Gamma\vdash_{\bf L}\varphi$ then there exists a finite subset $\Gamma_{0}$ of $\Gamma$ such that $\Gamma_{0}\vdash_{\bf L}\varphi$.
A very natural way to characterize (Tarskian) logics is by means of logical matrices, which formalize the idea of truth-tables:
Definition 2.1 (Logical matrices).
@@ 2 Preliminaries: Logics, Nmatrices and RNmatrices
Let $\Theta$ be a signature. A logical matrix over $\Theta$ is a triple $\mathcal{M}=\langle\mathcal{V},\mathcal{D},\mathcal{O}\rangle$ where
@@
    $\mathcal{V}$ is a non-empty set (of truth-values);
@@
    $\mathcal{D}\subseteq\mathcal{V}$ is a non-empty set (of designated values);
@@
    $\mathcal{O}$ is a function assigning, to each $\in\Theta_{n}$, a truth function $\tilde{\#}:\mathcal{V}^{n}\rightarrow\mathcal{V}$ (that is, $\mathcal{A}_{\mathcal{M}}=\langle\mathcal{V},\mathcal{O}\rangle$ is an algebra over $\Theta$).
Definition 2.2.
A valuation over a matrix $\mathcal{M}$ is a homomorphism $v:For(\Theta)\to\mathcal{A}_{\mathcal{M}}$. That is, for every $\in\Theta_{n}$ and for every $\varphi_{1},\ldots,\varphi_{n}\in For(\Theta)$: $v(\#(\varphi_{1},\ldots,\varphi_{n}))=\mathcal{O}(\#)(v(\varphi_{1}),\ldots,v(\varphi_{n}))$. The set of valuations over $\mathcal{M}$ is denoted by $Val(\mathcal{M})$. We say that $\varphi$ is a semantical consequence of $\Gamma$ w.r.t. $\mathcal{M}$, denoted by $\Gamma\models_{\mathcal{M}}\varphi$ if, for every $v\in Val(\mathcal{M})$: $v(\varphi)\in\mathcal{D}$ whenever $v(\gamma)\in\mathcal{D}$ for every $\gamma\in\Gamma$. We write $\models_{\mathcal{M}}\varphi$ instead of $\emptyset\models_{\mathcal{M}}\varphi$.
As mentioned in the Introduction, the formal notion of non-deterministic matrices (Nmatrices) was introduced by Avron and Lev in [2] as a way to overcome some Gödel-Dugundji’s uncharacterizability results. Finite-valued Nmatrices which are analytical (i.e., each partial valuation can be extended to a full valuation) provide a decision procedure for some logics which are uncharacterizable by a single finite-valued logic, such as some paraconsistent logics (see, for instance, [1]).
Definition 2.3 (Nmatrices, [2]).
Let $\Theta$ be a signature. A non-deterministic matrix (Nmatrix, for short) over $\Theta$ is a triple $\mathcal{M}=\langle\mathcal{V},\mathcal{D},\mathcal{O}\rangle$ where
  1. 1.
    $\mathcal{V}$ is a non-empty set (of truth-values);
  2. 2.
    $\mathcal{D}\subseteq\mathcal{V}$ is a non-empty set (of designated values);
  3. 3.
    $\mathcal{O}$ is a function assigning, to each $\in\Theta_{n}$, a (non-deterministic) truth function $\mathcal{O}(\#):\mathcal{V}^{n}\rightarrow(\wp(\mathcal{V}))\setminus\emptyset$ (that is, $\mathcal{A}_{\mathcal{M}}=\langle\mathcal{V},\mathcal{O}\rangle$ is a non-deterministic algebra over $\Theta$).
Definition 2.4 (Valuations on Nmatrices, [2]).
A valuation over an Nmatrix $\mathcal{M}$ is a function $v:For(\Theta)\to\mathcal{V}$ such that, for every $\in\Theta_{n}$ and for every $\varphi_{1},\ldots,\varphi_{n}\in For(\Theta)$, $v(\#(\varphi_{1},\ldots,\varphi_{n}))\in\mathcal{O}(\#)(v(\varphi_{1}),\ldots,v(\varphi_{n}))$. The set of valuations over $\mathcal{M}$ is denoted by $Val^{nd}(\mathcal{M})$. The notion of semantical consequence w.r.t. an Nmatrix $\mathcal{M}$, as well as the notation, will be analogous to the ones in Definition 2.2, but now w.r.t. valuations over the Nmatrix $\mathcal{M}$.
However, as mentioned before, certain logics are still uncharacterizable by finite-valued Nmatrices, such as da Costa’s paraconsistent logic $C_{1}$ (see [1]), modal logics KT and S4 (see [12]), or $\mathbf{IPL}$ (see Theorem 3.5 below). In these cases, finite-valued and decidable RNmatrices, which are analytical and co-analytical, can overcome this limitation.
Definition 2.5 (RNmatrices, [7]).
Let $\Theta$ be a signature. An RNmatrix over $\Theta$ is a pair $\mathcal{R}=\langle\mathcal{M},\mathcal{F}\rangle$ where $\mathcal{M}$ is an Nmatrix over $\Theta$ and $\mathcal{F}\subseteq Val^{nd}(\mathcal{M})$. The notion of semantical consequence w.r.t. an RNmatrix $\mathcal{R}=\langle\mathcal{M},\mathcal{F}\rangle$, as well as the notation, will be analogous to the ones in Definition 2.2, but now w.r.t. valuations in $\mathcal{F}$.
It is worth noting that the logic generated by an RNmatrix is indeed Tarskian.
$\neg^{S4}$
@@ 2 $\{0\}$ $\{1,2\}$ $\{1,2\}$
Example 2.6 (Grätz’s Nmatrix for S4).
Let $\Sigma_{b}=\{\{\neg,\Box\},\{\rightarrow\}\}$ and $\Sigma=\{\{\neg,\Box\},\{\rightarrow,\lor,\land\}\}$ be, respectively, the basic signature and the signature for modal logic S4. Given the set $\mathcal{P}$ of propositional variables, let $For(\Sigma_{b})$ and $For(\Sigma)$ be the corresponding propositional languages. Grätz’s Nmatrix over $\Sigma_{b}$ is the tuple $\mathcal{M}_{S4}=\langle\mathcal{V},\mathcal{D},\mathcal{O}\rangle$, where $\mathcal{V}=\{0,1,2\}$, $\mathcal{D}=\{1,2\}$, $\mathcal{O}=\{\neg^{S4},\Box^{S4},\rightarrow^{S4}\}$ and each $c\in\mathcal{O}$ is defined in Table 1. We will return to this Nmatrix in Section 5.
By setting $\varphi\lor\psi:=\neg\varphi\to\psi$ and $\varphi\land\psi:=\neg(\varphi\to\neg\psi)$, respectively, the multioperators $\lor^{S4}$ and $\land^{S4}$ are defined as in Table 1.
As proven in [12], no finite Nmatrix can characterize S4; in particular, $\mathcal{M}_{S4}$ cannot characterize this modal logic. In Section 5 we will analyze an RNmatrix based on $\mathcal{M}_{S4}$ which characterizes S4.
3 Gödel’s uncharacterizability theorem updated
As mentioned in the Introduction, Gödel proved in $1932$ that $\mathbf{IPL}$ cannot be characterized by a single finite-valued (deterministic) logical matrix. It is a natural question to determine whether $\mathbf{IPL}$ could be characterized by means of a single finite-valued non-deterministic matrix. As we shall show in the sequel, it is easy to adapt the proof by Gödel in [10] in order to see that the answer to the question above is negative.